OpenStudy (anonymous):

A storage chest is shaped like a cube. What is the volume of the storage chest when each side is x/4

OpenStudy (jacob902):

x3/64 cubic units

Mehek (mehek14):

Formula = L * W * H \(\dfrac{x}{4}*\dfrac{x}{4}*\dfrac{x}{4}=\dfrac{x^3}{64}\)